You invested ​$13000 in two accounts paying
2% and
6 %
annual​ interest, respectively.
If the total interest earned for the year was
$380, how much was invested at each​ rate?

Found 2 solutions by addingup, Alan3354:
Answer by addingup(3677)

Let the amount invested at 2% be x
Then, the amount invested at 6% is 13000 - x
.
0.02x + 0.06(13000 - x) = 380
0.02x + 780 - 0.06x = 380
-0.04x = -400
x = -400/-0.04 = divide, remember that -/- = + so you can ignore the - sign when you divide.
This will be the amount invested at 2%. Subtract the amount from 13000 to find out the amount invested at 6%.
.
Happy learning

Answer by Alan3354(69443)

"respectively" is superfluous in the statement.
--------
Some think it makes them appear intelligent to use it, but it's just the opposite.